OpenStudy (anonymous):

1. Which of the following values is the best approximation of the circumference of a circle with a diameter of 12 meters? (1 point) 6.0 m 18.9 m 36.0 m 37.7 m

OpenStudy (anonymous):

d=12, r=6 circumference = 2*pi*r =2*pi*6=12*3.14=37.7